6) How much energy would it take to change 55.0 g of water at-10.0°C to 20.0°C?
Answer #1

here,

mass , m = 55 g = 0.055 kg

specific heat of ice , Ci = 2090 J/kg degree C

latent heat of fusion , Lv = 334000 J/kg

specific heat of water , Cw = 4186 J/Kg degree C

the heat energy required , Q = m * ( Ci * ( 0 - (-10)) + Lf + Cw * ( 20 - 0))

Q = 0.055 * ( 2090 * 10 + 334000 + 4186 * ( 20))

Q = 2.41 * 10^4 J

the heat energy required is 2.41 * 10^4 J